OpenStudy (anonymous):

What is the value of the expression when a = 4 and b = 2? 2a^3 + 2b^2 i thought it was 48 three is the exponent and 2 is also the exponent ^

OpenStudy (anonymous):

2a^3+2b^2=2*4^3 + 2*2^2 = 2*64 + 2*4 = 128+8 = 136

OpenStudy (anonymous):

2*4*4*4+2*2*2 =128 + 8 = 136